Calculate the temporary hardness and total hardness of a sample of water containing magnesium bicarbonate, 73 mg/L; Calcium bicarbonate, 162 mg/L, magnesium chloride, 95 mg/L, Calcium sulphate, 136 mg/L


1
Expert's answer
2021-03-18T02:51:27-0400

calculation of CaCO3 equivalents:

Constituent Multiplication factorCaCO3 equivalent

Mg(HCO3)2= 73mg/L 100/146 73X100/146= 50mg/L 

Ca(HCO3)2= 162mg/L 100/162 162X100/162=100mg/L

MgCl2= 95mg/L 100/95 95X100/95= 100mg/L

CaSO4=136mg/L 100/136 136X100/136= 100mg/LT

Therefore, Temporary hardness of water due to magnesium biocarbonate and Calcium biocarbonate

=100 + 50 = 150

150mg/L or ppm.

Total hardness of water= 50+100+100+100=350 mg/L or ppm.